# Uni-3DAR: Unified 3D Generation and Understanding via Autoregression on Compressed Spatial Tokens

### v2

@@ 3.6 Polymer Property Prediction via Pretraining
Following prior research [78, 79], we use eight publicly available polymer property datasets (Egc, Egb, Eea, Ei, Xc, EPS, Nc, and Eat), obtained via density functional theory (DFT) calculations. Given that all tasks involve structure-level regression, we employ a robust evaluation strategy using 5-fold cross-validation with random splits, consistent with previous work [79]. Results are reported as the root mean squared error (RMSE), averaged across three different random seeds.
@@
Baseline methods include ChemBERTa [80], Uni-Mol [8], SML [78], PML [78], polyBERT [81], Transpolymer [82], and MMPolymer [79]. For pretraining, we represent homopolymers as specialized molecular structures using the star substitution strategy proposed in [79]. The model is pretrained using our masked next-token prediction strategy for 1 million steps with a batch size of 128. All other experimental details follow the settings previously described in the molecular property prediction experiments.
During fine-tuning, we adopt structure-level understanding strategy with masked next-token prediction auxiliary generative loss. Training is capped at 200 epochs. We thoroughly investigate various hyperparameter combinations by using three different batch sizes (32, 64, 128) and three learning rates (5e-4, 1e-4, 3e-4), creating nine unique configurations. Each configuration is tested by training models three times with different random seeds. We follow the same 5-fold split index align with [79], by averaging the best validation metrics in each fold. Subsequently, we present the mean performance along with the standard deviation across three seeds.
Results
The experimental results, summarized in Table 8, highlight Uni-3DAR’s competitive performance. Uni-3DAR ranks first in 4 out of 8 tasks and is within the top two in 7 out of 8 tasks. These findings underscore Uni-3DAR’s robust representational capabilities across molecular and homopolymer property prediction tasks.
@@ Results
ChemBERTa 0.539 $\pm$ 0.049 0.664 $\pm$ 0.079 0.350 $\pm$ 0.036 0.485 $\pm$ 0.086 18.711 $\pm$ 1.396 0.603 $\pm$ 0.083 0.140 $\pm$ 0.010 0.219 $\pm$ 0.056
Uni-Mol 0.489 $\pm$ 0.028 0.531 $\pm$ 0.055 0.332 $\pm$ 0.027 0.407 $\pm$ 0.080 17.414 $\pm$ 1.581 0.536 $\pm$ 0.053 0.095 $\pm$ 0.013 0.084 $\pm$ 0.034
SML 0.489 $\pm$ 0.056 0.547 $\pm$ 0.110 0.313 $\pm$ 0.016 0.432 $\pm$ 0.060 18.981 $\pm$ 1.258 0.576 $\pm$ 0.020 0.102 $\pm$ 0.010 0.062 $\pm$ 0.014
PLM 0.459 $\pm$ 0.036 0.528 $\pm$ 0.081 0.322 $\pm$ 0.037 0.444 $\pm$ 0.062 19.181 $\pm$ 1.308 0.576 $\pm$ 0.060 0.100 $\pm$ 0.010 0.050 $\pm$ 0.010
polyBERT 0.553 $\pm$ 0.011 0.759 $\pm$ 0.042 0.363 $\pm$ 0.037 0.526 $\pm$ 0.068 18.437 $\pm$ 0.560 0.618 $\pm$ 0.049 0.113 $\pm$ 0.003 0.172 $\pm$ 0.016
Transpolymer 0.453 $\pm$ 0.007 0.576 $\pm$ 0.021 0.326 $\pm$ 0.040 0.397 $\pm$ 0.061 17.740 $\pm$ 0.732 0.547 $\pm$ 0.051 0.096 $\pm$ 0.016 0.147 $\pm$ 0.093
MMPolymer 0.431 $\pm$ 0.017 0.503 $\pm$ 0.038 0.286 $\pm$ 0.029 0.390 $\pm$ 0.057 16.814 $\pm$ 0.867 0.511 $\pm$ 0.035 0.087 $\pm$ 0.010 0.061 $\pm$ 0.016
Uni-3DAR 0.426 $\pm$ 0.022 0.498 $\pm$ 0.048 0.291 $\pm$ 0.022 0.396 $\pm$ 0.072 17.16 $\pm$ 1.498 0.487 $\pm$ 0.034 0.087 $\pm$ 0.011 0.066 $\pm$ 0.031
3.7 Mutual Benefits of Generation and Understanding Tasks
@@
Atom Sta(%$\uparrow$ Mol Sta(%$\uparrow$ Valid(%$\uparrow$ V $\times$ U(%$\uparrow$
@@ 3.8 Ablation Study
Atom Sta(%$\uparrow$ Mol Sta(%$\uparrow$ Valid(%$\uparrow$ V $\times$ U(%$\uparrow$ Tokens Cost $\downarrow$
@@
    Masked Next-Token Prediction significantly enhances generation performance. In experiment No.2, we followed previous work [20] that merely appends the position of the next token to the current token, without using our proposed masked next-token prediction. Comparing experiments No.1 and No.2 clearly demonstrates that our proposed masked next-token prediction substantially outperforms this baseline approach.
@@
    Token-level diffusion loss yields comparable performance to the autoregressive head but with lower efficiency. Our default generation head uses a simple autoregressive head (refer to Alg. 1) to sequentially predict atom types and in-cell positions. We examined whether employing a more powerful head, such as the token-level diffusion loss from MAR [23], could further enhance performance. Experiment No.6, utilizing the diffusion head, achieved similar results but required more computational time. Therefore, we opt for the simpler, more efficient autoregressive head by default.
@@
    Combining atom-based autoregressive and diffusion losses without spatial tokenization is insufficient. Recent works have explored improving atom-based autoregressive generation through token-level diffusion losses [45]. We tested this approach by adding a token-level diffusion loss to experiment No.5, resulting in experiment No.7. Although No.7 performed slightly better than No.5, it remained significantly inferior to the proposed Uni-3DAR. This underscores that comprehensive spatial information, as provided by our tokenization strategy, is crucial, mere integration of diffusion-based methods into atom-based autoregressive model, without spatial tokenization, cannot achieve substantial performance improvements.
3.9 Inference Speed
We benchmarked Uni-3DAR against the diffusion-based generative model GeoLDM [12] on QM9 generation task (Sec.3.1) by evaluating the throughput (i.e., the number of molecules generated per second). Model throughput was evaluated across a range of batch sizes, with all experiments conducted on a single Nvidia 4090 GPU. As shown in Fig.4, Uni-3DAR consistently outperforms the diffusion-based approach in sampling efficiency, achieving significantly reduced generation times across all tested settings. In particular, at larger batch sizes, Uni-3DAR is approximately 21.8x faster than GeoLDM, and even at a small batch size of 64, it remains about 7.5x faster. Additionally, we assessed the inference overhead introduced by masked next-token prediction. Thanks to our optimizations (Sec. 2.4), we find that masked next-token prediction incurs only a 15% to 30% slowdown. Given its substantial performance gains, this additional cost is well justified.
@@ 3.9 Inference Speed
4 Related Work
Octree and Hierarchical Autoregressive Models
The coarse-to-fine hierarchical structure is widely used in 3D vision [83, 84, 85, 86, 87, 20, 88, 89]. Among these works, [20] is most similar to our Uni-3DAR, as it also employs autoregressive generation using an octree. However, our method differs in three key aspects: (1) instead of relying on deep tree-level generation for fine details, we add an extra layer of fine-grained tokens to avoid excessively deep trees; (2) rather than compressing nodes via convolutional layers, we represent a compressed subtree with a single token; and (3) to handle dynamic token positions, while [20] appends the next position to the current token, we adopt a masked next-token prediction strategy. These innovations make Uni-3DAR more efficient and effective than [20]. Recently, some image generative models have adopted a coarse-to-fine, level-by-level generation approach, such as VAR [90]. Although the high-level idea appears similar, our motivation is distinct: Uni-3DAR is designed to avoid the inefficiencies of a full-size cubic grid, whereas VAR uses more tokens to boost performance. Moreover, Uni-3DAR remains within the next-token prediction framework, while VAR employs next-scale prediction to parallelize token generation at each level.
Microscopic 3D Structure Modeling
Most previous generative models for microscopic 3D structures employ diffusion-based approaches [35, 91, 11, 12, 38] to generate atomic positions from noise. However, diffusion models have two major limitations. First, they require the number of atoms to be predetermined. Second, atom types are sampled from a categorical distribution, for which a proper score function is not well defined. Some studies have explored grid-based generation [18] to address these issues, but using a full-size 3D grid is computationally prohibitive. Other works have investigated autoregressive models for 3D molecules [14, 15, 45], but these models generate molecules atom by atom, requiring a predefined sequence order. Furthermore, during generation, the position of the next atom is unknown and must be predicted along with its type, making the process significantly more challenging and affecting overall performance.
For microscopic 3D structure understanding, prior studies primarily leverage SE(3)-invariant or equivariant models [92, 93]. Additionally, unsupervised pretraining is widely used to mitigate the scarcity of labeled data [75, 94, 74, 13, 8]. These models typically follow a BERT-style pretraining framework [95], where some atoms are masked, their 3D positions are perturbed, and the model is trained to recover the ground truth. While highly effective for understanding tasks, most of these models cannot be directly applied to generation.
Some recent efforts have attempted to unify generation and understanding for microscopic data. However, most focus solely on sequence data (e.g., 1D SMILES, nucleotide sequences, or textual descriptions) and directly apply autoregressive language models [96, 97, 98, 99]. While these models are straightforward, they lack essential 3D structural information, limiting their performance and applicability. Recent studies have also explored diffusion-based approaches. For example, UniGEM [36] demonstrated that a two-phase, multi-task training strategy can improve performance for both tasks. This approach combines diffusion loss with a prediction task applied during later diffusion steps. In summary, while previous work has made progress in bridging generation and understanding, Uni-3DAR is the first autoregressive framework to unify both tasks for 3D microscopic structure modeling.
Macroscopic 3D Structure Modeling
In machine learning, macroscopic 3D structure modeling encompasses the understanding and generation of everyday objects [100, 101], scenes [102], CAD models [103, 104, 105], avatars [106], and more. Similar to microscopic 3D structures, macroscopic 3D structures lack a unified representation format. Commonly used 3D representations include voxels [107], point clouds [108], polygon meshes [109], implicit functions [85], and 3D Gaussian Splatting [110]. Recent methods, such as 3DShape2VecSet [111], Michelangelo [112], CLAY [88], and Dora [113], encode 3D shapes into compressed, compact latent codes, substantially improving representation efficiency. Specifically, 3DShape2VecSet introduces a transformer-based 3D VAE that encodes uniformly sampled surface points into a compact latent code, upon which CLAY further builds using the latent Diffusion Transformer (DiT) [114]. Dora advances this line of work by improving both the surface sampling strategy and the 3D VAE architecture. Previous literature also explored autoregressive modeling for macroscopic 3D structures. For example, Polygen [115] and MeshGPT [116] generate mesh faces sequentially from lowest to highest on the vertical axis, corresponding to the point-based tokenization strategy as discussed in Sec. 1, suffering from the same challenges in dynamic token positions. All of these approaches, however, differ fundamentally from the proposed Uni-3DAR. Our unified framework can be applied directly to macroscopic datasets or combined with state-of-the-art representations for potentially even stronger performance.
Another category of 3D structure generation methods, known as optimization-based approaches [117, 118, 119, 120], leverages text-to-image generative models and refines 3D representations by distilling information from 2D images [120]. Unlike true 3D generation, these methods primarily perform 3D reconstruction, making them fundamentally distinct from the previously mentioned 3D generation techniques and Uni-3DAR.
5 Conclusion
In this paper, we introduced Uni-3DAR, a unified autoregressive framework for 3D structural generation and understanding. By leveraging a novel coarse-to-fine tokenization method based on an octree, along with 2-Level Subtree Compression and Masked Next-Token Prediction, Uni-3DAR efficiently captures the full spatial context while avoiding cubic token growth. Extensive experiments across a range of tasks on microscopic structures demonstrate that our approach consistently outperforms or matches existing methods while being significantly more efficient.
For future work, we aim to extend Uni-3DAR to a broader range of 3D structures, particularly macroscopic ones. Additionally, jointly training a large-scale Uni-3DAR foundation model on diverse data sources and tasks, as well as integrating it with other modalities, presents an exciting research direction.
Author Contributions
Acknowledgments
Appendix A Illustration of the Generated Examples

### v3

@@ C.7 Polymer Property Prediction via Pretraining
Following prior research [Zhang et al., 2023b, Wang et al., 2024], we use eight publicly available polymer property datasets (Egc, Egb, Eea, Ei, Xc, EPS, Nc, and Eat), obtained via density functional theory (DFT) calculations. Given that all tasks involve structure-level regression, we employ a robust evaluation strategy using 5-fold cross-validation with random splits, consistent with previous work [Wang et al., 2024]. Results are reported as the root mean squared error (RMSE), averaged across three different random seeds.
@@
Baseline methods include ChemBERTa [Chithrananda et al., 2020], Uni-Mol [Zhou et al., 2023b], SML [Zhang et al., 2023b], PML [Zhang et al., 2023b], polyBERT [Kuenneth and Ramprasad, 2023], Transpolymer [Xu et al., 2023a], and MMPolymer [Wang et al., 2024]. For pretraining, we represent homopolymers as specialized molecular structures using the star substitution strategy proposed in [Wang et al., 2024]. The model is pretrained using our masked next-token prediction strategy for 1 million steps with a batch size of 128. All other experimental details follow the settings previously described in the molecular property prediction experiments.
During fine-tuning, we adopt structure-level understanding strategy with masked next-token prediction auxiliary generative loss. Training is capped at 200 epochs. We thoroughly investigate various hyperparameter combinations by using three different batch sizes (32, 64, 128) and three learning rates (5e-4, 1e-4, 3e-4), creating nine unique configurations. Each configuration is tested by training models three times with different random seeds. We follow the same 5-fold split index align with [Wang et al., 2024], by averaging the best validation metrics in each fold. Subsequently, we present the mean performance along with the standard deviation across three seeds.
@@
ChemBERTa [Chithrananda et al., 2020]
0.539
$\pm$ 0.049
0.664
$\pm$ 0.079
0.350
$\pm$ 0.036
0.485
$\pm$ 0.086
18.711
$\pm$ 1.396
0.603
$\pm$ 0.083
0.140
$\pm$ 0.010
0.219
$\pm$ 0.056
Uni-Mol [Zhou et al., 2023b]
0.489
$\pm$ 0.028
0.531
$\pm$ 0.055
0.332
$\pm$ 0.027
0.407
$\pm$ 0.080
17.414
$\pm$ 1.581
0.536
$\pm$ 0.053
0.095
$\pm$ 0.013
0.084
$\pm$ 0.034
SML [Zhang et al., 2023b]
0.489
$\pm$ 0.056
0.547
$\pm$ 0.110
0.313
$\pm$ 0.016
0.432
$\pm$ 0.060
18.981
$\pm$ 1.258
0.576
$\pm$ 0.020
0.102
$\pm$ 0.010
0.062
$\pm$ 0.014
PLM [Zhang et al., 2023b]
0.459
$\pm$ 0.036
0.528
$\pm$ 0.081
0.322
$\pm$ 0.037
0.444
$\pm$ 0.062
19.181
$\pm$ 1.308
0.576
$\pm$ 0.060
0.100
$\pm$ 0.010
0.050
$\pm$ 0.010
polyBERT [Kuenneth and Ramprasad, 2023]
0.553
$\pm$ 0.011
0.759
$\pm$ 0.042
0.363
$\pm$ 0.037
0.526
$\pm$ 0.068
18.437
$\pm$ 0.560
0.618
$\pm$ 0.049
0.113
$\pm$ 0.003
0.172
$\pm$ 0.016
Transpolymer [Xu et al., 2023a]
0.453
$\pm$ 0.007
0.576
$\pm$ 0.021
0.326
$\pm$ 0.040
0.397
$\pm$ 0.061
17.740
$\pm$ 0.732
0.547
$\pm$ 0.051
0.096
$\pm$ 0.016
0.147
$\pm$ 0.093
MMPolymer [Wang et al., 2024]
0.431
$\pm$ 0.017
0.503
$\pm$ 0.038
0.286
$\pm$ 0.029
0.390
$\pm$ 0.057
16.814
$\pm$ 0.867
0.511
$\pm$ 0.035
0.087
$\pm$ 0.010
0.061
$\pm$ 0.016
Uni-3DAR
0.426
$\pm$ 0.022
0.498
$\pm$ 0.048
0.291
$\pm$ 0.022
0.396
$\pm$ 0.072
17.16
$\pm$ 1.498
0.487
$\pm$ 0.034
0.087
$\pm$ 0.011
0.066
$\pm$ 0.031
Appendix D More Experiments
D.1 Mutual Benefits of Generation and Understanding Tasks
@@
[…]
@@ D.2 Ablation Study
[…]
@@
    Masked Next-Token Prediction significantly enhances generation performance. In experiment No.2, we followed previous work [Ibing et al., 2023] that merely appends the position of the next token to the current token, without using our proposed masked next-token prediction. Comparing experiments No.1 and No.2 clearly demonstrates that our proposed masked next-token prediction substantially outperforms this baseline approach.
@@
    Token-level diffusion loss yields comparable performance to the autoregressive head but with lower efficiency. Our default generation head uses a simple autoregressive head (refer to Alg. 1) to sequentially predict atom types and in-cell positions. We examined whether employing a more powerful head, such as the token-level diffusion loss from MAR [Li et al., 2024b], could further enhance performance. Experiment No.6, utilizing the diffusion head, achieved similar results but required more computational time. Therefore, we opt for the simpler, more efficient autoregressive head by default.
@@
    Combining atom-based autoregressive and diffusion losses without spatial tokenization is insufficient. Recent works have explored improving atom-based autoregressive generation through token-level diffusion losses [Zhang et al., 2025]. We tested this approach by adding a token-level diffusion loss to experiment No.5, resulting in experiment No.7. Although No.7 performed slightly better than No.5, it remained significantly inferior to the proposed Uni-3DAR. This underscores that comprehensive spatial information, as provided by our tokenization strategy, is crucial, mere integration of diffusion-based methods into atom-based autoregressive model, without spatial tokenization, cannot achieve substantial performance improvements.
D.3 Inference Speed
We benchmarked Uni-3DAR against the diffusion-based generative model GeoLDM [Xu et al., 2023b] on QM9 generation task (Sec.3.1) by evaluating the throughput (i.e., the number of molecules generated per second). Model throughput was evaluated across a range of batch sizes, with all experiments conducted on a single Nvidia 4090 GPU. As shown in Fig.4, Uni-3DAR consistently outperforms the diffusion-based approach in sampling efficiency, achieving significantly reduced generation times across all tested settings. In particular, at larger batch sizes, Uni-3DAR is approximately 21.8x faster than GeoLDM, and even at a small batch size of 64, it remains about 7.5x faster. Additionally, we assessed the inference overhead introduced by masked next-token prediction. Thanks to our optimizations (Sec. B), we find that masked next-token prediction incurs only a 15% to 30% slowdown. Given its substantial performance gains, this additional cost is well justified.
@@ D.3 Inference Speed
Appendix E Illustration of the Generated Examples
@@
Figure SI-4: Macroscopic 3D object generation samples of ShapeNet dataset.